OpenStudy (anonymous):

i really need help please, i will medal and fan Pies are ordered from the Pizza Palace according to their diameter. Large:sixteen inches Medium:fourteen inches Small:twelve inches How many more square inches are in the large pizza than in the small pizza? twelve point five six square inches forty-seven point one zero square inches eighty-seven point nine two square inches three hundred fifty-one point six eight square inches

OpenStudy (jessiegonzales):

Area of Circle = pi r^2 large pizza has 16" diameter or 8" radius A = pi (8)^2 A = 201.06 square inches small pizza has 12" diamerter or 6" radius A = pi (6)^2 A = 113.1 subtract area of large pizza from small pizza 201.06 - 113.10 = 87.96 square inches

OpenStudy (anonymous):

i think it would be 87.92 @jessiegonzales

OpenStudy (jessiegonzales):

Sorry my mistake typed the wrong number!!! I showed step by step though no medal? THanks @hns941 for correcting me :)

OpenStudy (anonymous):

sorry lol i thought i clicked it @jessiegonzales

OpenStudy (jessiegonzales):

THankyou!!! Any other questions? Let me know :)

OpenStudy (anonymous):

i have another open one i will tag you :)